# Probing the axion-photon coupling with space-based gravitational wave detectors

### v1

@@ Abstract
We propose a simple modification of space-based gravitational wave (GW) detector optical benches which would enable the measurement of vacuum birefringence of light induced by axion dark matter through its coupling to electromagnetism. While marginally affecting the sensitivity to GW by a factor $\sqrt{2}$, we show that such an adjustment would make future detectors such as LISA, TianQin, Taiji and Big-Bang Observer the most sensitive experiments at low axion masses.
@@ II Vacuum birefringence
where $R$ is the Ricci scalar, $g_{\mu\nu}$ is the spacetime metric tensor (with ($+$)(-\>+\>+\>+)( - + + + ) signature) with determinant $g$ and $\kappa=8\pi G/c^{4}$ is the Einstein gravitational constant. At cosmological scales, when $m_{a}c^{2}/\hbar\gg H$, with $H$ the Hubble constant, the axion field equation derived from Eq. (1) admits plane wave solutions, i.e.
@@
where $\rho_{\mathrm{DM}}=0.4$ GeV/cm3 is the local DM energy density McMillan (2011), $\vec{v}_{\mathrm{DM}}=v_{\mathrm{DM}}\hat{e}_{v}\sim 10^{-3}\>c\hat{e}_{v}$ is the mean velocity of the Sun in the galactic halo and
@@
In addition to the action Eq. (1), we add the Lagrangian describing the interaction in vacuum between the axion $a$ and electromagnetism (EM) through the coupling $g_{a\gamma}$ Sikivie (2021)
@@ III Optical benches
In this section, we will be interested in the polarization of light inside the detectors OB and in between the S/C. As a reminder, all the experiments under consideration in this paper will consist of one (or more) constellation of three S/C, in the following noted $1,2,3$. Each S/C is hosting two OB (labeled $A$ and $B$) with one laser per OB. These two laser beams are sent to the two other S/C and GW signals will be detected by interferometric measurement between those optical signals (see Fig. 1). For our purposes, we are interested in the polarization of these beams inside the S/C and also in between the S/C.
@@ III.2 Requirements for GW detection
  • When leaving one S/C, a small portion of the local beam will be undesirably reflected on it, come back inside the OB and create an additional noise in the interferometer [^1] Assuming the frequency of the local beam to be $f_{\ell}$, the distance beam has a slightly different frequency $f_{\ell}+\delta f$, due to standard Doppler from S/C differential velocity, and frequency planning, see Heinzel et al. (2024). The reflected component of the local beam could interfere with itself but would only create a DC component. However, it could also interfere with the distant beam, and create a spurious beatnote at $\delta f$. In addition, the motion of the reflecting component at the edge of the S/C would induce a time dependent phase on the reflected beam, which would mimic a signal from the GW, and greatly alter the measurement.. We require this reflected component to be rejected by the polarized beam splitter, such that it does not enter the interferometer.
It can be shown easily that the current OB design from Fig. 2 fulfills those requirements. In the full constellation (with the three S/C) with two OB in each S/C, six individual interference measurements are performed at each time. Two different linear polarizations travel in space : vertical polarization travels in one direction (say clockwise) while horizontal polarization travels in the other direction (say counterclockwise) (see Fig. 1).
@@ IV Sensitivity of space-based GW observatory to the axion-photon coupling
In the previous section, we have demonstrated that a minor change in the OB design from space-baed GW observatories can make them sensitive to the coupling between an axion and electromagnetism. Let us now assess quantitatively the sensitivity of these space missions to $g_{a\gamma}$.
@@ IV.1 Single-arm Doppler effect
where $\omega_{j}/2\pi\approx\omega_{0}/2\pi=\nu_{0}$ is the laser frequency, supposed to be the same inside all OB.
@@ IV.2 Sagnac $\alpha$ TDI combination
In the time domain, the second generation Sagnac $\alpha$ combination is defined as Hartwig et al. (2023)
$\alpha(t)$ $=-\left(1-D_{1231}\right)\left[y_{13}+D_{13}y_{32}+D_{132}y_{21}-\right.\,$
$\left.\left(y_{12}+D_{12}y_{23}+D_{123}y_{31}\right)\right]\,,$ (13a)
@@
with $\Phi^{\prime\prime}=\Phi^{\prime}-5\omega_{a}L/2c=\Phi-3\omega_{a}L/c$.
The one-sided power spectral density (PSD) of this signal reads
@@
The modification of the OB suggested in Fig. 3 implies a decrease in the amplitude of light (or equivalently in the number of photons $N_{\gamma}$) that reaches each interferometer by a factor $2$ because light gets partially absorbed by the right polarizer. Since the OMS noise includes a contribution from shot noise whose PSD $\propto 1/N_{\gamma}$, we will conservatively add a factor $2$ to the whole OMS contribution [^4] The sensitivity to GW strain scaling as the square root of the TDI noise PSD, this implies a factor $\sqrt{2}$ loss of sensitivity on $h$., such that the Sagnac $\alpha$ noise PSD reads Hartwig et al. (2023) [^5] In Hartwig et al. (2023), only the first generation noise PSD are presented (see their Eq. (B2a)). In order to compute the second generation noise PSD for the Sagnac $\alpha$ combination, one needs to add an additional factor $(2\sin(3\omega L/2c))^{2}$.
@@ IV.3 Expected sensitivity
Galactic DM is not purely monochromatic, it is a superposition of plane waves which oscillate at slightly different frequencies (see e.g. Derevianko (2018); Foster et al. (2018)). Therefore, the field acquires a characteristic coherence time $\tau(f)\sim 10^{6}/f$ Derevianko (2018). Eq. (17a) is valid when $T_{\mathrm{obs}}\ll\tau(f)$, i.e when the field can effectively be described as a monochromatic plane wave. Note that a correction factor to the sensitivity arises due to the stochastic nature of the amplitude of the field Centers et al. (2021) which induces a loss in signal of $\sim 1.5$ for a SNR 1. When the integration time is much longer than the coherence time of the field, i.e $T_{\mathrm{obs}}\gg\tau(f)$, the signal searched for is no longer coherent, i.e. it should be modeled as a sum of several stochastic harmonics, see Foster et al. (2018), and therefore Eq. (14) is not valid anymore. Another method to analyse the data is to cut the dataset in fragments with duration smaller than $\tau(f)$ and search for a coherent signal in each of these blocks of data. In such a case, the experimental sensitivity to the coupling is reduced and becomes Budker et al. (2014)
@@ V Discussion
As we have shown it in Fig. 5, operating this new OB would allow the various space-based GW detectors under consideration in this paper to reach unconstrained regions of the $g_{a\gamma}-m_{a}$ parameter space, improving current best constraints on $g_{a\gamma}$ by as much as 5 orders of magnitude over several orders of magnitude of axion masses. In particular, such detection techniques would be complementary to futuristic versions of GW ground-based detectors, optimized for the axion-photon coupling searches Nagano et al. (2019), which are the most sensitive for axion masses between $\sim 10^{-16}$ and $10^{-11}$ eV (due to the smaller baseline).
@@ Acknowledgements.
This work was supported by the Programme National GRAM of CNRS/INSU with INP and IN2P3 cofunded by CNES.

### v2

@@ Abstract
We propose a simple modification of space-based gravitational wave (GW) detector optical benches which would enable the measurement of vacuum birefringence of light induced by axion dark matter through its coupling to electromagnetism. Specifically, we propose to change a half-wave plate by a circular polarizer. While marginally affecting the sensitivity to GW by a factor $\sqrt{2}$, we show that such an adjustment would make future detectors such as LISA, TianQin, Taiji and Big-Bang Observer the most sensitive experiments at low axion masses.
@@ II Vacuum birefringence
where $R$ is the Ricci scalar, $g_{\mu\nu}$ is the spacetime metric tensor (with ($+$)(-\>+\>+\>+)( - + + + ) signature) with determinant $g$ and $\kappa=8\pi G/c^{4}$ is the Einstein gravitational constant. At cosmological scales, when $m_{a}c^{2}/\hbar\gg H$, with $H$ the Hubble constant, the axion field equation derived from Eq. (1) admits plane wave solutions, i.e. [^1] In reality, galactic DM is a stochastic superposition of plane waves (see e.g. Foster et al. (2018)) that has a finite coherence time. Therefore, the observables of interest in this paper (Eqs. (8c), (12) and (13b)) will not be purely monochromatic. For the frequencies of interest here, this stochastic feature will not change significantly the sensitivity estimates.
@@
where $\rho_{\mathrm{DM}}=0.4$ GeV/cm3 is the local DM energy density McMillan (2011) [^2] The actual numerical value of $\rho_{\mathrm{DM}}$ depends on the model used to describe the velocity distribution of stars around the galaxy and can vary from $\sim 0.3$ GeV/cm3 to $\sim 0.7$ GeV/cm3, see Cirelli et al. (2024); Evans et al. (2019). This is why most DM sensitivity plots (like the one we show in Fig. 5) assume a specific value of $\rho_{\mathrm{DM}}$ for consistent comparison., $\vec{v}_{\mathrm{DM}}=v_{\mathrm{DM}}\hat{e}_{v}\sim 10^{-3}\>c\hat{e}_{v}$ is the mean velocity of the Sun in the galactic halo and
@@
In addition to the action Eq. (1), we add the Lagrangian describing the interaction in vacuum between the axion $a$ and EM through the coupling $g_{a\gamma}$ Sikivie (2021)
@@ III Optical benches
In this section, we will be interested in the polarization of light inside the detectors OB and in free space, between the S/C. As a reminder, all the experiments under consideration in this paper will consist of one (or more) constellation of three S/C, in the following noted $1,2,3$. Each S/C is hosting two OB (labeled $A$ and $B$) with one laser per OB. These two laser beams are sent to the two other S/C and GW signals will be detected by interferometric measurement between those optical signals (see Fig. 1). For our purposes, we are interested in the polarization of these beams inside the S/C and also in between the S/C.
@@ III.2 Requirements for GW detection
  • When leaving one S/C, a small portion of the local beam will be undesirably reflected on it, coming back inside the OB to create an additional noise in the interferometer [^3] Assuming the frequency of the local beam to be $f_{\ell}$, the distance beam has a slightly different frequency $f_{\ell}+\delta f$, due to standard Doppler from S/C differential velocity, and frequency planning, see Heinzel et al. (2024). The reflected component of the local beam could interfere with itself but would only create a DC component. However, it could also interfere with the distant beam, and create a spurious beatnote at $\delta f$. In addition, the motion of the reflecting component at the edge of the S/C would induce a time dependent phase on the reflected beam, which would mimic a signal from the GW, and greatly alter the measurement.. We require this reflected component to be rejected by the polarized beam splitter, such that it does not enter the interferometer.
It can be shown easily that the current OB design from Fig. 2 fulfills those requirements [^4] There are likely several additional effects that need to be carefully studied before an implementation of the modification that we propose (e.g. additional stray light from the polarizer itself, angular or translational motion of the beam induced by the polarizer, phase fluctuations induced by the polarizer, etc…). Evaluating them in detail is beyond the scope of this paper, but we note that those effects are similar to the “usual” effects already studied and under control for other optical components, like the $\lambda/2$ wave plates in the original design. We therefore assume that suitable solutions can be found, and restrain this paper to introducing the principle and estimating the corresponding sensitivity.. In the full constellation (with the three S/C) with two OB in each S/C, six individual interference measurements are performed at each time. Two different linear polarizations travel in space : vertical polarization travels in one direction (say clockwise) while horizontal polarization travels in the other direction (say counterclockwise) (see Fig. 1).
@@ IV Sensitivity of space-based GW observatory to the axion-photon coupling
In the previous section, we have demonstrated that a minor change in the OB design from space-based GW observatories can make them sensitive to the coupling between an axion and electromagnetism. Let us now assess quantitatively the sensitivity of these space missions to $g_{a\gamma}$.
@@ IV.1 Single-arm Doppler effect
where $\omega_{j}/2\pi\approx\omega_{0}/2\pi=\nu_{0}$ is the laser frequency, supposed to be the same inside all OB. As it can be noticed from Eq. (12), the Doppler shift induced by the axion-photon coupling is unpolarized in the sense that it does not depend on e.g. the polarization of the incoming wave (as it is the case for the GW, see e.g. Cornish and Rubbo (2003)) or the propagation direction of the incoming wave (as it is the case for other DM couplings, see e.g. Yu et al. (2023); Morisaki and Suyama (2019)). Therefore, this signal differs significantly from other signals expected to be detected in the space missions of interest in this paper.
@@ IV.2 Sagnac $\alpha$ TDI combination
In the time domain, the second generation Sagnac $\alpha$ combination is defined as (see the definition of $C_{1}^{12}$ in Table 1 from Hartwig and Muratore (2022))
$\alpha(t)$ $=\left(1-D_{1321}\right)\left(y_{12}+D_{12}y_{23}+D_{123}y_{31}\right)$
$-\left(1-D_{1231}\right)\left(y_{13}+D_{13}y_{32}+D_{132}y_{21}\right)\,,$ (13a)
@@
with $\Phi^{\prime\prime}=\Phi^{\prime}-5\omega_{a}L/2c=\Phi-3\omega_{a}L/c$. While the signal has been derived assuming constant and equal arm-lengths, real S/C orbits will induce unequal arm-lengths with time variation at the relative level of the percent (see Fig. 5 from Martens and Joffre (2021)). This will impact the amplitude of the signal from Eq. (13b) (and subsequently the sensitivity on $g_{a\gamma}$) at the percent level at maximum for axion masses below $10^{-15}$ eV. At higher masses, only the exact location of the resonances will be impacted at the percent level. The one-sided power spectral density (PSD) of this signal reads
@@
The modification of the OB suggested in Fig. 3 implies a decrease in the amplitude of light (or equivalently in the number of photons $N_{\gamma}$) that reaches each interferometer by a factor $2$ because light gets partially absorbed by the right polarizer (a linear polarization being a superposition of left and right states, the right polarizer transmits only half of the light intensity from the linear input). Since the OMS noise includes a contribution from shot noise whose PSD $\propto 1/N_{\gamma}$, we will conservatively add a factor $2$ to the whole OMS contribution [^7] The sensitivity to GW strain scaling as the square root of the TDI noise PSD, this implies a factor $\sqrt{2}$ loss of sensitivity on $h$., such that the Sagnac $\alpha$ noise PSD reads Hartwig et al. (2023) [^8] In Hartwig et al. (2023), only the first generation noise PSD are presented (see their Eq. (B2a)). In order to compute the second generation noise PSD for the Sagnac $\alpha$ combination, one needs to add an additional factor $(2\sin(3\omega L/2c))^{2}$.
@@ IV.3 Expected sensitivity
Galactic DM is not purely monochromatic, it is a superposition of plane waves which oscillate at slightly different frequencies (see e.g. Derevianko (2018); Foster et al. (2018)). Therefore, the field acquires a characteristic coherence time $\tau(f)\sim 10^{6}/f$ Derevianko (2018). Eq. (17a) is valid when $T_{\mathrm{obs}}\ll\tau(f)$, i.e when the field can effectively be described as a monochromatic plane wave. Note that a correction factor to the sensitivity arises due to the stochastic nature of the amplitude of the field Centers et al. (2021) which induces a loss in signal of $\sim 1.5$ for a SNR of 1. When the integration time is much longer than the coherence time of the field, i.e $T_{\mathrm{obs}}\gg\tau(f)$, the signal searched for is no longer coherent, i.e. it should be modeled as a sum of several stochastic harmonics, see Foster et al. (2018), and therefore Eq. (14) is not valid anymore. Another method to analyse the data is to cut the dataset in fragments with duration smaller than $\tau(f)$ and search for a coherent signal in each of these blocks of data. In such a case, the experimental sensitivity to the coupling is reduced and becomes Budker et al. (2014)
@@ V Discussion
As we have shown it in Fig. 5, operating this new OB would allow the various space-based GW detectors under consideration in this paper to reach unconstrained regions of the $g_{a\gamma}-m_{a}$ parameter space, improving current best constraints on $g_{a\gamma}$ by as much as 5 orders of magnitude over several orders of magnitude of axion masses. In particular, such detection techniques would be complementary to futuristic versions of GW ground-based detectors, optimized for the axion-photon coupling searches Nagano et al. (2019), which are the most sensitive for axion masses between $\sim 10^{-16}$ and $10^{-11}$ eV (due to the smaller baseline). Note that the modification that we propose would not be optimal for ground based detectors, because they use Fabry-Perot cavities. Indeed, assuming circular states, the polarization of photons change at each reflection inside the cavity. Therefore, the phase accumulated over two reflections will in general cancel (because of the sign difference in Eq. (8a)), except if the round trip time corresponds exactly to a period of oscillation of the axion field.
@@ Acknowledgements.
This work was supported by the Programme National GRAM of CNRS/INSU with INP and IN2P3 cofunded by CNES. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. J.G. is funded by the grant CNS2023-143767. Grant CNS2023-143767 funded by MICIU/AEI/10.13039/501100011033 and by European Union NextGenerationEU/PRTR.